### v1

@@ 6 Discussion
We observe that the vacuum flow in the causal diamond has a special feature: the observers are following trajectories of constant acceleration — see footnote 58. For general geometric flows in $2d$ CFTs, this is not the case, so there is no direct interpretation of the observer’s experience in terms of a thermal bath at temperature $\frac{1}{\beta_{\mathrm{phys}}}$. One interesting situation is where $\nabla_{\mu}\xi^{\mu}$ is small and the observer’s experience is approximately thermal. [^62] In $d=2$, this means $|c^{\prime}(u)+b^{\prime}(v)|\ll 1$. This condition holds, for example, for the thermal-state flow (36) in the Rindler wedge whenever $u$ is large and negative, and $v$ is large and positive (compared to $\beta$).
@@
Furthermore, even when the observer experiences a thermal bath with $\beta=\beta_{\rm phys}$, our description of the entropy and energy, e.g., in eqs. (140) and (161), is only heuristic, since the resulting Unruh temperatures do not describe a CFT plasma with a varying temperature which is in local thermal equilibrium. In order to have local thermal equilibrium, the typical thermal wavelength should be smaller than the length scale over which the temperature is varying. By contrast, for example, the vacuum modular flow in the Rindler wedge on the $t=0$ slice has typical thermal wavelength $\lambda\simeq\beta=2\pi x$, which is of the same order of magnitude of the length scale over which the temperature varies, given by $x$.
@@
To close then, we return to the original question: In what sense does $\beta=2\pi\norm{\xi}$ correspond to a temperature? A precise notion of local temperatures has been provided in Arias:2016nip ; Arias:2017dda . In these works, one excites the vacuum with a unitary operator localized at a spacetime point $\boldsymbol{w}$, yielding $\Delta S_{EE}=0$ yet $\Delta\langle h\rangle\neq 0$. In particular, $S_{rel}=\Delta\langle h\rangle$, and if the modular Hamiltonian has the form (5), one has $S_{rel}\approx 2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})$, where $P_{\mu}$ is the expectation value of the total energy-momentum of the excitation. The relative entropy $S_{rel}=S_{rel}(\rho_{\Psi_{1}}^{\cal V}||\rho_{\Psi_{2}}^{\cal V})$ measures how difficult it is to operationally distinguish $\Psi_{1}$ and $\Psi_{2}$ using measurements in $\mathcal{V}$. That is, the probability of confounding the two states after $N$ measurements is $p\sim e^{-S_{rel}N}$. If $\Psi_{1}$ differs from $\Psi_{2}$ by a localized unitary excitation such that $S_{rel}\approx 2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})$, then we find the excitation in the quantum fluctuations of $\Psi_{1}$ with the same probability as in a thermal state with “covariant temperature” $2\pi\xi^{\mu}(\mathbf{w})$, see Arias:2016nip ; Arias:2017dda . As an example, consider the Rindler wedge; setting $\boldsymbol{w}=(0,a)$ in $(t,x)$ coordinates, one finds $2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})=2\pi T_{00}\xi^{t}(\boldsymbol{w})=2\pi aE$, retrieving the Rindler temperature $T=\frac{1}{2\pi a}$.
@@
There is therefore a precise sense in which $2\pi\xi^{\mu}$ becomes a covariant temperature and $\beta=2\pi\norm{\xi}$ is the inverse temperature of an observer along the modular flow. While $\beta_{\mathrm{phys}}$ is a physical temperature measured by a detector, $\beta=2\pi\norm{\xi}$ corresponds to an information-theoretic temperature related to the distinguishability and entanglement of states.
@@ Acknowledgements.
We thank Michele Bianchessi, Horacio Casini, Laurent Freidel, Samuel Goldman, Tom Hartman, Ted Jacobson, Marc Klinger, Rob Leigh, Caroline Lima, Tassos Petkou, Maria Preciado-Rivas, Kasia Rejzner, Antony Speranza, Eirini Telali, Marija Tomašević, Erik Tonni, Themistocles Zikopoulos for useful discussions and feedback. We especially thank Jon Sorce for key discussions and feedback on the draft. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. JC acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number 514036]. RCM is also supported in part by a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada, and by funding from the BMO Financial Group. FC is partially supported by University of Milan through the Scholarship for Master’s Degree Thesis Preparation Abroad, and by Professor Florian Girelli through the Graduate Research Studentship at University of Waterloo.
@@ Appendix A Flows under conformal maps
$\langle\Psi|\sigma_{s+2\pi i}^{\Psi,\mathcal{V}}({\sf a})\,{\sf b}|\Psi\rangle=\langle\Omega|\sigma_{s+2\pi i}^{\Omega,\varphi^{-1}(\mathcal{V})}(\mathsf{A})\,\mathsf{B}|\Omega\rangle=\langle\Omega|\mathsf{B}\,\sigma_{s}^{\Omega,\varphi^{-1}(\mathcal{V})}(\mathsf{A})|\Omega\rangle=\langle\Psi|{\sf b}\,\sigma_{s}^{\Psi,\mathcal{V}}({\sf a})|\Psi\rangle\,,$ (194)
@@ Appendix F Multiple intervals
In this appendix, we apply our entropy formula eq. (140) to an example not considered in the main text. In particular, we show that it reproduces the correct entanglement entropy even when the modular Hamiltonian contains additional nonlocal terms.
In particular, in free fermion theory in $d=2$, the modular Hamiltonian for a pair of interals contains both local and non-local terms Casini:2009vk . In the case of two symmetric intervals $x\in(-R,-r)\cup(r,R)$ with $r\in(0,R)$, on the $t=0$ slice, the local part reads Casini:2016fgb
@@
Since this term is not of the form (5), the vacuum modular flow with respect to two intervals is non-geometric. One expects our entropy formula (140) to apply only for states with geometric modular flows. So, let us imagine there exists another state whose modular Hamiltonian with respect to the two intervals is given simply by $h_{\mathrm{loc}}$, and calculate the resulting entanglement entropy. We will be surprised to recover the correct entanglement entropy for the standard vacuum state.
First, note $h_{\mathrm{loc}}$ can be generalized beyond the case of symmetric intervals in the $t=0$ slice. The more general operator reads Arias:2016nip
@@
By (140), the two intervals have an entanglement entropy of
@@
This result straightforwardly generalizes to the case of more than two intervals, and matches the entanglement entropy calculated in Casini:2009sr for $p$ subregions. [^72] Note that in Casini:2009sr , the points $X_{i}$ and $Y_{i}$ lie on the same $t=0$ slice, whereas this is not required here. The interesting point here is that we recovered the correct result for the entanglement entropy using only the local term (278) in the modular Hamiltonian, suggesting the nonlocal term does not actually contribute.
Note, while we have recovered the entanglement entropy in the vacuum state by instead considering a modified state whose modular Hamiltonian does not include the nonlocal term eq. (277), we do not expect to reproduce the mutual information in the vacuum state. This is because the nonlocal term encodes entanglement between the two regions (rather than between the joint subregion and its complement), and thus we expect a vanishing mutual information between the two regions in the modified state. On the other hand, the mutual information between these two regions in the vacuum is nonvanishing.

### v2

@@ 6 Discussion
As an aside, recall that the vacuum flow in the causal diamond has a special feature: the observers are following trajectories of constant acceleration. For general geometric flows in $2d$ CFTs, this is not the case, and there is no direct interpretation of the observer’s experience in terms of a thermal bath at temperature $\frac{1}{\beta_{\mathrm{phys}}}$. One interesting situation is where $\nabla_{\mu}\xi^{\mu}$ is small and the observer’s experience is approximately thermal. [^66] In $d=2$, this means $|c^{\prime}(u)+b^{\prime}(v)|\ll 1$. This condition holds, for example, for the thermal-state flow (36) in the Rindler wedge whenever $u$ is large and negative, and $v$ is large and positive (compared to $\beta$).
@@
Further, even when the observer experiences a thermal bath with $\beta=\beta_{\rm phys}$, our description of the entropy and energy, e.g., in eqs. (140) and (161), is only heuristic, since the resulting Unruh temperatures do not describe a CFT plasma with a varying temperature which is in local thermal equilibrium. In order to have local thermal equilibrium, the typical thermal wavelength should be smaller than the length scale over which the temperature is varying. By contrast, for example, the vacuum modular flow in the Rindler wedge on the $t=0$ slice has typical thermal wavelength $\lambda\simeq\beta=2\pi x$, which is of the same order of magnitude of the length scale over which the temperature varies, given by $x$.
Nevertheless, our thermodynamics-inspired entropy formula (140) for states with geometric modular flow $\xi$ is on good footing. Recall $\xi$ is a future-directed conformal Killing vector in the subregion of interest, $\mathcal{V}$. Using a singular Weyl transformation, $\xi\big{|}_{\mathcal{V}}$ be transformed into an ordinary Killing field which is future-directed and timelike everywhere; this happened, for example, in section 5.1, when conformally mapping the double cone to a hyperbolic spacetime. In the new conformal frame where $\xi$ is Killing, the modular temperature equals the physical temperature, and this temperature is varying over very large length scales (or not at all, in the double-cone example). Hence, the thermodynamic entropy formula (138) applies. This validates the entanglement entropy formula (140) because by footnote 47, eq. (140) is Weyl-invariant, and hence it is also applicable for the original setting where $\xi$ is only conformally Killing.
@@
To close, we revisit the original question: In what sense does $\beta=2\pi\norm{\xi}$ correspond to a temperature? A precise notion of local temperatures has been provided in Arias:2016nip ; Arias:2017dda . In these works, one excites the vacuum with a unitary operator localized at a spacetime point $\boldsymbol{w}$, yielding $\Delta S_{EE}=0$ yet $\Delta\langle h\rangle\neq 0$. In particular, $S_{rel}=\Delta\langle h\rangle$, and if the modular Hamiltonian has the form (5), one has $S_{rel}\approx 2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})$, where $P_{\mu}$ is the expectation value of the total energy-momentum of the excitation. The relative entropy $S_{rel}=S_{rel}(\rho_{\Psi_{1}}^{\cal V}||\rho_{\Psi_{2}}^{\cal V})$ measures how difficult it is to operationally distinguish $\Psi_{1}$ and $\Psi_{2}$ using measurements in $\mathcal{V}$. That is, the probability of confounding the two states after $N$ measurements is $p\sim e^{-S_{rel}N}$. If $\Psi_{1}$ differs from $\Psi_{2}$ by a localized unitary excitation such that $S_{rel}\approx 2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})$, then we find the excitation in the quantum fluctuations of $\Psi_{1}$ with the same probability as in a thermal state with “covariant temperature” $2\pi\xi^{\mu}(\mathbf{w})$, see Arias:2016nip ; Arias:2017dda . As an example, consider the Rindler wedge; setting $\boldsymbol{w}=(0,a)$ in $(t,x)$ coordinates, one finds $2\pi P_{\mu}\xi^{\mu}(\boldsymbol{w})=2\pi T_{00}\xi^{t}(\boldsymbol{w})=2\pi aE$, retrieving the Rindler result $T=\frac{1}{2\pi a}$.
@@
There is therefore a precise sense in which $2\pi\xi^{\mu}$ can be interpreted as a covariant temperature and $\beta=2\pi\norm{\xi}$ is the inverse temperature of an observer along the modular flow. While $\beta_{\mathrm{phys}}$ is a physical temperature measured by a detector, $\beta=2\pi\norm{\xi}$ corresponds to an information-theoretic temperature related to the distinguishability and entanglement of states.
@@ Acknowledgements.
We thank Michele Bianchessi, Horacio Casini, Laurent Freidel, Samuel Goldman, Tom Hartman, Ted Jacobson, Marc Klinger, Rob Leigh, Caroline Lima, Tassos Petkou, Maria Preciado-Rivas, Pranav Pulakkat, Kasia Rejzner, Mykola Semenyakin, Antony Speranza, Eirini Telali, Marija Tomašević, Erik Tonni, Themistocles Zikopoulos for useful discussions and feedback. We especially thank Jon Sorce for key discussions and feedback on the draft. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. JC acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number 514036]. RCM is also supported in part by a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada, and by funding from the BMO Financial Group. FC is partially supported by University of Milan through the Scholarship for Master’s Degree Thesis Preparation Abroad, and by Professor Florian Girelli through the Graduate Research Studentship at University of Waterloo.
@@ Appendix A Flows under conformal maps
$\langle\Psi|\sigma_{s-2\pi i}^{\Psi,\mathcal{V}}({\sf a})\,{\sf b}|\Psi\rangle=\langle\Omega|\sigma_{s-2\pi i}^{\Omega,\varphi^{-1}(\mathcal{V})}(\mathsf{A})\,\mathsf{B}|\Omega\rangle=\langle\Omega|\mathsf{B}\,\sigma_{s}^{\Omega,\varphi^{-1}(\mathcal{V})}(\mathsf{A})|\Omega\rangle=\langle\Psi|{\sf b}\,\sigma_{s}^{\Psi,\mathcal{V}}({\sf a})|\Psi\rangle\,,$ (196)
@@ Appendix F Multiple intervals
In this appendix, we apply our entropy formula eq. (140) to an example not considered in the main text: the case of two intervals in $d=2$ free fermion theory.
In this case, the modular Hamiltonian contains both local and non-local terms Casini:2009vk . In the case of two symmetric intervals $x\in(-R,-r)\cup(r,R)$ with $r\in(0,R)$, on the $t=0$ slice, the local part reads Casini:2016fgb
@@
Since this term is not of the form (5), the vacuum modular flow with respect to two intervals is non-geometric. One expects our entropy formula (140) to apply only for states with geometric modular flows. Fortunately, from the discussion in section 6, we know there exists another state |$\ket{\Psi_{p}}$ whose modular Hamiltonian with respect to the two intervals is given simply by $h_{\mathrm{loc}}$, and we can apply the entanglement entropy formula to this state instead. We will show that |$\ket{\Psi_{p}}$ and |$\ket{\Omega}$ have the same entanglement entropy.
To do so, first note that $h_{\mathrm{loc}}$ can be generalized beyond the case of symmetric intervals in the $t=0$ slice. The more general operator reads Arias:2016nip
@@
By eq. (140), the two intervals in the state |$\ket{\Psi_{p}}$ have an entanglement entropy of
@@
This result straightforwardly generalizes to the case of more than two intervals, and matches the vacuum state entanglement entropy calculated in Casini:2009sr for $p$ subregions. [^76] Note that in Casini:2009sr , the points $X_{i}$ and $Y_{i}$ lie on the same $t=0$ slice, whereas this is not required here. The interesting point here is that we recovered the correct result for the vacuum entanglement entropy using only the local term (280) in the modular Hamiltonian, suggesting the nonlocal term does not actually contribute. Of course, we do not expect to recover the correct mutual information, since |$\ket{\Psi_{p}}$ has no entanglement between the two subregions, and in particular the mutual information vanishes. On the other hand, the mutual information between two subregions in the vacuum state is nonvanishing.